06natrod52378:

3x-5=34 I got {X=13} Is that correct? Just checking to make sure my hw is done correctly

unknowngirly:

im tryna think

unknowngirly:

yep ur right

06natrod52378:

ok ty

unknowngirly:

becuase multiply 3 times 13 and u get 39 then take away 5 and u get that

unknowngirly:

@06natrod52378 wrote:
ok ty
np

umm:

Correct. Problem:3x−5=34 Step by step: Add 5 to both sides \[3x−5+5=34+5\] \[3x=39\] Divide both sides by 3 \[\frac{ 3x }{ 3 }=\frac{ 39 }{ 3 }\] Therefore \[x=13\]

06natrod52378:

Is this the correct way to show work?

1 attachment
unknowngirly:

look's right to me 😊

06natrod52378:

can i just post my whole homework page? Its inished just need to make sure it all correct

unknowngirly:

i think u can idk lol it's a hw site

06natrod52378:

1 attachment
umm:

@06natrod52378 wrote:
can i just post my whole homework page? Its inished just need to make sure it all correct
Is it a test, or homework?

06natrod52378:

Hw

umm:

Then yes, you're fine.

06natrod52378:

1 attachment
unknowngirly:

The first is one is right

06natrod52378:

ik that lol u already said it was